OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (1, –3) and has a slope of 2? 2x – y = 5 2x – y = –5 2x + y = 5 2x + y = –5

OpenStudy (ivettef365):

which one you think is correct

OpenStudy (anonymous):

i don't know yet.

OpenStudy (ivettef365):

use formula y=mx + b, to find value of b so you have points (x,y) -> (1,-3) insert in formula and m= slope = 2 -3 = (2)(1) + b so what is the value of b ??

OpenStudy (anonymous):

i don't know how to find that out.

OpenStudy (anonymous):

someone please help.

OpenStudy (ivettef365):

well you have -3 = 2 +b -2 -2 ================= -5 = b so you have b = -5

OpenStudy (anonymous):

so it's either B or D

OpenStudy (ivettef365):

now you know that the equation is y = 2x -5

OpenStudy (anonymous):

is it B?

OpenStudy (ivettef365):

no, because let see B how would look like when I transfer you need to isolate y B) 2x -y = -5 -y = -2x - 5 ---- --- --- -1 -1 -1 y = 2x +5 not good

OpenStudy (anonymous):

ohh so it's D?

OpenStudy (ivettef365):

Lets try D) 2x + y = -5 y = -2x -5 not good because the slope shows -2

OpenStudy (anonymous):

what the..... okay A?

OpenStudy (anonymous):

Ok wait, The equation of line in slope point form is given as; \[y-y _{1}=m(x-x _{1})\] \[y-(-3)=2(x-1) \rightarrow y+3=2x-2 \rightarrow y=2x-2 -3 \rightarrow y=2x-5\] I.e. 2x-y=5 is the required eq of the line

OpenStudy (anonymous):

yayyy it's A

OpenStudy (ivettef365):

ok, let try A A) 2x -y = 5 -y = -2x +5 ---- --- --- -1 -1 -1 y = 2x - 5 GOOD